A real estate salesperson bought promotional calendars and date books to give to her customers at the end of the year. The calendars cost $0.75 each, and the date books cost $0.50 each. She ordered a total of 500 promotional items and spent $300. How many of each item did she order?
Answer by Cromlix(4381)

Calendars = x
Books = y
x + y = 500........1
75x + 50y = 30000......2
Multiply (1) by 50
50x + 50y = 25000.....3
75x + 50y = 30000.....2
Subtract (3) from (2)
25x = 5000
Divide both sides by 25
x = 200
Substitute x = 200 in
Eq(1)
x + y = 500
200 + y = 500
y = 500 - 200
y = 300
Therefore, the salesperson ordered
200 calendars and 300 date books.
Hope this helps.
:-)